please show all steps

Omar has an investment worth 168,187 dollars. The investment will make a special payment of X dollars to Omar in 7 month(s) and the investment also will make regular, fixed monthly payments of 1,780 dollars to Omar forever. The expected return for the investment is 1.15 percent per month and the first regular, fixed monthly payment of 1,780 dollars will be made to Omar in 1 month. What is X, the amount of the special payment that will be made to Omar in 7 month(s)?

An investment, which is worth 30,300 dollars and has an expected return of 7.42 percent, is expected to pay fixed annual cash flows forever with the next annual cash flow expected in 1 year. What is the present value of the annual cash flow that is expected in 7 years from today?
Answer #1

Present value of fixed monthly payments

= 1,780/1.15%

= $154,782.61

Present value of payment to be received at the end of 7 months = $168,187-154,782.61

= $13,404.39

Amount of X = 13,404.39(1+0.0115)7

= $14,521.39

Let Annual cash flow be x

30,300 = x/7.42%

X = $2,248.26

Present value of cash flows expected from 7 years from today = 2,248.26*PVF(7.42%, 7 years)

= 2,248.26*0.606

= $1,362.45